OpenStudy (anonymous):

Write a function that is NOT linear with a domain of -3 < x, x∈ |R and range of 0 < y, y∈ |R.

OpenStudy (mr.math):

\(y=\sqrt{x+3}\), the domain is \(x\ge -3\) and the range is \(y\ge 0\).